a particle moves in a ciorcle of radius 21/22 mwith constant speed 1m/s.find:


a)magnitude of av velocity


b)magnitude of av acceleration in 2 secs.

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
12 years ago

Dear Student

Particle will cover a distance of 2m in 2sec.

circumference of circle= 2*22/7*21/22 = 6m

in 6m angular displacement is 3600. so in 2m it will be 1200

join centre to the initial and final position of the particle. it will be a triangle with 1200 angle at the centre and other two angle equal to 300 each.  length of cord will give the displacement.

if rad is R then

displacement= 2Rcos300= 2*21/22*√3 /2 = 21√3/2

avg vel =disp/time= 21√3/4

 

let at t=0 v1= i (i is unit vector along x axis) and particle is moving clockwise

at t=2 , v2= -sin300 i + -cos300 j= -i/2 - √3 /2 j

v2-v1= -3i/2 - √3 /2 j

avg acc= (-3i/2 - √3 /2 j)/2

 

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

 

 

AKASH GOYAL

AskiitiansExpert-IIT Delhi